# On Function-Correcting Codes in the Lee Metric

### v2

@@ 8 Conclusion
In this paper, we extended the study of function-correction codes for the Lee metric by introducing function-correcting Lee metric codes (FCLMCs) over the residue ring $\mathbb{Z}_{m}$. FCLMCs require less redundancy compared to classical Lee metric codes. In this encoding, we only need to correct those codewords whose associated function values differ, rather than correcting all erroneous codewords. We derived upper and lower bounds on the optimal redundancy of FCLMCs for general functions utilizing the smallest length irregular Lee distance codes. Moreover, we defined locally Lee functions, and using general results, we obtained bounds on the optimal redundancy for locally Lee functions. We also showed that the lower bound on optimal redundancy can be attained in the case of locally binary Lee functions. We further explored the locality of specific functions, namely the Lee weight function and the Lee weight distribution function, and established bounds on the optimal redundancy for these functions. Our results generalize the corresponding bounds over $\mathbb{Z}_{2}$ given in [5] and over $\mathbb{Z}_{4}$ given in [6]. Additionally, we established a Plotkin-like bound for linear function-correcting codes in the Lee metric. Potential directions for future work include extending the analysis of FCLMCs to a specific class of functions of interest and examining the tightness of the bounds obtained in this paper.
Acknowledgment
We thank Hareesh K., Rashid Ummer N.T., and B. Sundar Rajan for pointing out a mistake in Lemma 4.1. We corrected the lemma and added Section 7 in this updated version.

### v3

@@ 8 Conclusion
In this paper, we extended the study of function-correction codes for the Lee metric by introducing function-correcting Lee metric codes (FCLMCs) over the residue ring $\mathbb{Z}_{m}$. FCLMCs require less redundancy compared to classical Lee metric codes. In this encoding, we only need to distinguish those codewords whose associated function values differ, rather than correcting all erroneous codewords. We derived upper and lower bounds on the optimal redundancy of FCLMCs for general functions utilizing the smallest length irregular distance codes in the Lee metric. Moreover, we defined locally bounded functions in the Lee metric, and using general results, we obtained bounds on the optimal redundancy for locally bounded functions. We also showed that the lower bound on optimal redundancy can be attained in the case of locally binary functions. We further explored the locality of specific functions, namely the Lee weight function and the Lee weight distribution function, and established bounds on the optimal redundancy. Moreover, we provided explicit constructions for these functions. Our results generalize the corresponding bounds over $\mathbb{Z}_{2}$ given in [6] and over $\mathbb{Z}_{4}$ given in [7]. Additionally, we explored linear FCLMCs and established a Plotkin-like bound for linear function-correcting codes in the Lee metric.
A natural and nontrivial continuation of this work is the investigation and construction of coding schemes that provide simultaneous protection against errors in both the function outputs and the underlying message data. Developing such codes, capable of jointly safeguarding data and function values in the Lee metric, constitutes a challenging and promising direction for future research, with potential impact on error-resilient computation, storage, and communication systems. Recently, in [11], the authors studied such code constructions in the Hamming metric.